3) Prove by induction that for all natural numbers n:


(a) 1 + 3 + 5 + · · · + (2n − 1) = n2




1
Expert's answer
2021-05-26T03:52:14-0400

Let "P(n)" be the proposition that the sum "1+3+5+...+(2n-1)" is "n^2" for all natural numbers "n"


"1+3+5+...+(2n-1)=n^2"

Basic step

"P(1)" is true, because "1=1^2."


Inductive step

For the inductive hypothesis we assume that "P(k)" holds for an arbitrary

natural number "k." That is, we assume that


"1+3+5+...+(2k-1)=k^2"

Under this assumption, it must be shown that "P(k+1)" is true, namely, that


"1+3+5+...+(2(k+1)-1)=(k+1)^2"

When we add "(2(k+1)-1)" to both sides of the equation in "P(k)," we obtain


"1+3+5+...+(2k-1)+(2(k+1)-1))"

"=k^2+(2(k+1)-1)"

"1+3+5+...+(2k-1)+(2(k+1)-1))="

"=k^2+2k+1"


"1+3+5+...+(2k-1)+(2(k+1)-1))="

"=(k+1)^2"


This last equation shows that "P(k+1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all natural numbers "n". That is, we have proven that


"1+3+5+...+(2n-1)=n^2"

for all natural numbers "n."